OpenStudy (anonymous):

In parallelogram DEFG, DH= x+3, HF=3y, GH=3x-3, and HE=5Y+4. Find the values of x and y.

OpenStudy (anonymous):

where is the H belonging to...........

OpenStudy (anonymous):

H is the perpendicular bisector.

OpenStudy (anonymous):

x+3=3y 3x-3=5Y+4 x=3y-3________put this in second equation 3(3y-3)-3=5y+4 3y-9-3=5y+4 2y=16 y=8 x=21

OpenStudy (anonymous):

http://www.mathopenref.com/parallelogram.html